A Gaussian surface is a closed surface in three-dimensional space through which the electric flux is calculated. Given a spherical Gaussian surface that has a radius of 0.5 meters and encloses 30 electronsdetermine the value of the electric field at a distance equal to 0.6 meters from the center of the surface. (nN/C) 


1
Expert's answer
2021-05-09T13:03:43-0400
"E=\frac{nq}{4\pi\epsilon_0 r^2}\\
E=\frac{30(1.6\cdot10^{-19})}{4\pi(8.85\cdot10^{-12})(0.6)^2}=120\frac{nN}{C}"